# Diffusion Model Patching via Mixture-of-Prompts

### v1

@@ 4 Experiments
Pre-trained (iter: 7M)
   DiT-XL/2 2.29
Further Training (iter: 20K)
   + Fine-tuning 5.04(+2.75)
   + Prompt tuning 2.77(+0.48)
   + DMP 2.25(-0.04)
Table 2: Patching the pre-trained DiT models [43] with DMP. We set two baselines for comparison: (1) conventional fine-tuning to update the model parameters. (2) naive prompt tuning [29] (equivalent to Eq. 4). Note that we use the same dataset as in the pre-training. $\uparrow$/$\downarrow$: The higher/lower the better. Image Resolution is 256$\times$256.
4.1 Effectiveness of DMP
In Table 2, we present a comparison of DMP against two further training baselines – (1) traditional fine-tuning and (2) naive prompt tuning [29] (equivalent to Eq. 4) – across three datasets for various image generation tasks. We employ pre-trained DiT models [43] that have already reached full convergence as our backbone. To ensure that the observed enhancement is not due to cross-dataset knowledge transfer, we further train the models using the same dataset used for pre-training. As expected, fine-tuning does not provide additional improvements to an already converged model in all datasets and can even result in overfitting. Naive prompt tuning also fails to improve performance in all datasets and can instead lead to a decrease in performance. DMP enhances the FID across all datasets with only 20$\sim$40K iterations, enabling the model to generate images of superior quality. This indicates that the performance gains achieved by DMP are not merely a result of increasing parameters, but rather from its novel mixture-of-prompts strategy. This strategy effectively patches diffusion models to operate slightly differently at each denoising step.
[…]
Pre-trained (iter: 250K)
   DiT-L/2 6.26
@@ 4.1 Effectiveness of DMP
   + Fine-tuning 6.80(+0.54)
   + Prompt-tuning 6.51(+0.25)
   + DMP 5.61(-0.65)
@@
Table 3: Commparison of further training techniques on FFHQ [25].
@@
In Table 3, further training of a fully converged DiT-L/2 model reveals interesting dynamics. First, fine-tuning fails to increase performance beyond its already converged state and even tends to overfit, leading to performance degradation. We also found that prompt tuning, which uses a small number of extra parameters, actually harms performance, possibly because these extra parameters act as noise in the model’s input space. In contrast, DMP, which also uses the same set of parameters as prompt tuning, significantly boosts performance. The key difference between them lies in the use of a gating function: prompts are shared across all timesteps, while DMP activates prompts differently for each timestep. This distinction allows DMP, with a fixed number of parameters, to scale across thousands of timesteps by creating mixtures-of-prompts. By patching stage-specificity into the pre-trained diffusion model, DMP achieves a 10.38% FID gain in just 50K iterations.
4.2 Design Choices
Prompt depth.
To investigate the impact of the number of blocks in which prompts are inserted, we conduct an ablation study using the DiT-B/2 model, which comprises 12 DiT blocks. We evaluate the performance differences when applying mixtures-of-prompts at various depths in Fig. 4: only at the first block, up to half of the blocks, and across all blocks. Regardless of the depth, performance consistently improves compared to the baseline (FID=6.27) with no prompts. Our findings indicate a positive correlation between prompt depth and performance, with better results achieved using mixture-of-prompts across more blocks. The prompts selected for each block are illustrated in Fig. 5.
Gating architecture.
Our DMP framework incorporates a dynamic gating mechanism to select mixture-of-prompts specialized to each denoising step (or noise level). We compare the impact of two gating architectures in Table 4(a): linear gating vs. attention gating. The linear gating utilizes a single linear layer, taking a timestep embedding as input to produce a soft weighting mask for each learnable prompt. On the other hand, attention gating utilizes an attention layer [59], treating learnable prompts as a query and timestep embeddings as key and value, resulting in weighted prompts that are directly added to the latent vectors. Upon comparing the two gating architectures, we found that simple linear gating achieves better FID (5.87) compared to more sophisticated attention gating (6.41), which is actually worse than the baseline (6.27). As a result, we adopt linear gating as our default setting.
@@ 4.2 Design Choices
Table 4: DMP ablations. DiT-B/2 [43] pre-trained on FFHQ 256$\times$256 [25] is further trained for 30K iterations with DMP (Baseline FID = 6.27).
Gating type.
@@
In Table 4(b), we evaluate two design choices for creating mixture-of-prompts: hard selection vs. soft selection. With hard selection, we choose the top-$k$ prompts based on the gating function output probabilities, using them directly with a weight of 1. This explicitly separates the combination of prompts for each timestep. We set $k$-192 out of 256 prompts (75% of total prompts). On the other hand, soft selection uses all prompts but assigns different weights to each. Soft selection leads to further improvement, whereas hard selection results in performance that is actually worse than the base pre-trained model. Therefore, we set the soft selection as our default setting.
Prompt selection.
@@
By default, DMP inserts learnable prompts into the input space of every block of the diffusion model. Two choices arise in this context: (1) uniform: gating function $\mathcal{G}$ in Eq. 5 receives only the timestep embedding ${\bm{t}}$ as input and applies output weights to the prompts at every block, thus prompt selection is consistent across all blocks. (2) distinct: $\mathcal{G}$ processes not only ${\bm{t}}$ but also current block depth $i$ as inputs, generating different weights for each block. As shown in Table 4(c), using distinct prompt selections leads to enhanced performance. Therefore, we incorporate both the timestep embedding and current block depth information into the gating function, enabling the use of distinct prompt combinations for each block depth as our default setting.
Prompt position.
@@
Previous prompt tuning approaches [31, 29, 22, 71, 61] typically prepend learnable prompts to image tokens. In our work, we directly add prompts element-wise to image tokens, thereby maintaining the input sequence length. Table 4(d) ablates different choices for inserting prompts into the input space and their impact on performance. We compare two methods: prepend vs. add. For “add”, we use 256 prompts to match the number of image tokens, and for “prepend”, we utilize 50 prompts. Although ideally, “prepend” should also use 256 tokens for a fair comparison, we empirically found that prepending 256 prompt tokens causes diffusion models to diverge, even when both methods are initialized with zero equally. Hence, we limit “prepend” to 50 tokens. The results show that while “add” method improves performance with a stable optimization process, achieving a 5.87 FID compared to the baseline of 6.27 FID. On the other hand, the “prepend” leads to a drop in performance, with an FID of 6.79. Additionally, “add” has the advantage of not increasing computation overhead. Based on these findings, we set “add” as our default for stable optimization.
Prompt balancing.

### v3

@@ Introduction
Pre-trained (iter: 1.5M)
   Stable Diffusion v1.5 47.18
@@
   + Fine-tuning 52.99(+5.81)
   + Prompt tuning 35.93(-11.25)
   + DMP 35.44(-${}_{\textbf{(-11.74)}}$
Table 2: Evaluating pre-trained diffusion models with different further training methods. Importantly, we use the same dataset as in the pre-training for further training. We set two baselines for comparison: (1) full fine-tuning to update the entire model parameters. (2) naive prompt tuning (Lester, Al-Rfou, and Constant 2021) (equivalent to Eq. 2). $\downarrow$: The lower the better.
Experiments
We evaluate the effectiveness of DMP on various image generation tasks using already converged pre-trained diffusion models. Unlike conventional fine-tuning or prompt tuning, the original dataset from the pre-training phase is used for further training. We evaluated image quality using FID (Heusel et al. 2017) score, which measures the distance between feature representations of generated and real images using an Inception-v3 model (Szegedy et al. 2016).  [^4] Implementation details are in Appendix.
Datasets & Tasks. We used three datasets for our experiments: (1) FFHQ (Karras, Laine, and Aila 2019) (for unconditional image generation) contains 70,000 training images of human faces. (2) MS-COCO (Lin et al. 2014) (for text-to-image generation) includes 82,783 training images and 40,504 validation images, each annotated with 5 descriptive captions. (3) Laion5B (Schuhmann et al. 2022) (for Stable Diffusion) consists of 5.85B image-text pairs, which is known to be used to train Stable Diffusion (Rombach et al. 2022).
Comparative Study
Effectiveness of DMP.
In Table 2, we compare DMP against two further training baselines – (1) full fine-tuning and (2) naive prompt tuning (Lester, Al-Rfou, and Constant 2021) (equivalent to Eq. 2) – across various datasets for unconditional/conditional image generation tasks. We employ pre-trained DiT models (Peebles and Xie 2022) that have already reached full convergence as our backbone. To ensure that the observed enhancement is not due to cross-dataset knowledge transfer, we further train the models using the same dataset used for pre-training. As expected, fine-tuning does not provide additional improvements to an already converged model in all datasets and even result in overfitting. Naive prompt tuning also fails to improve performance in almost datasets and instead lead to a decrease in performance. DMP enhances the FID across all datasets with only 30$\sim$50K iterations, enabling the model to generate images of superior quality. This indicates that the performance gains achieved by DMP are not merely a result of increasing parameters, but rather from its novel mixture-of-prompts strategy. This strategy effectively patches diffusion models to operate slightly differently at each denoising step. Moreover, the significant improvement on Stable Diffusion v1.5 (Rombach et al. 2022) with Laion5B (Schuhmann et al. 2022) demonstrates DMP’s expandability to different architectures and resolutions.
FFHQ $256\times 256$ Further training iterations
DiT-L/2 (iter: 250K) +0K +10K +20K +30K +40K +50K
+ Fine-tuning 6.26 6.32 6.53 6.64 6.73 6.80(+0.54)
+ Prompt tuning 6.26 6.30 6.36 6.40 6.42 6.51(+0.25)
+ DMP 6.26 5.88 5.72 5.67 5.64 5.61(-0.65)
@@
Table 3: Comparison of further training techniques across iterations on FFHQ 256$\times$256.
Effects across training iterations.
@@
In Table 3, further training of a DiT-L/2 model reveals interesting dynamics. First, fine-tuning fails to increase performance beyond its already converged state and even tends to overfit, leading to performance degradation. We also found that prompt tuning, which uses a small number of extra parameters, actually harms performance, possibly because these extra parameters act as noise in the model’s input space. In contrast, DMP, which also uses the set of parameters as prompt tuning, significantly boosts performance. The key difference between them lies in the use of a gating function: prompts are shared across all timesteps, while DMP activates prompts differently for each timestep. This distinction allows DMP, with a fixed number of parameters, to scale across thousands of timesteps by creating mixtures-of-prompts. By patching stage-specificity into the pre-trained diffusion model, DMP achieves a 10.38% FID gain in just 50K iterations.
@@
Table 4: DMP ablations. DiT-B/2 (Peebles and Xie 2022) pre-trained on FFHQ 256$\times$256 (Karras, Laine, and Aila 2019) is further trained for 30K iterations with DMP (Baseline FID = 6.27). $\downarrow$: The lower the better.
Design Choices
Prompt depth. To investigate the impact of the number of blocks in which prompts are inserted, we conduct an ablation study using the DiT-B/2 model, which comprises 12 DiT blocks. We evaluate the performance differences when applying mixtures-of-prompts at various depths in Fig. 4: only at the first block, up to half of the blocks, and across all blocks. Regardless of the depth, performance consistently improves compared to the baseline with no prompts (FID=6.27). Our findings indicate a positive correlation between prompt depth and performance, with better results achieved using mixture-of-prompts across more blocks. The prompts selected for each block are illustrated in Fig. 5.
Gating architecture. Our DMP framework incorporates a dynamic gating mechanism to select mixture-of-prompts. We compare the impact of two gating architectures in Table 4(a): linear gating vs. attention gating. The linear gating utilizes a single linear layer, taking a timestep embedding as input to produce a weighting mask for each learnable prompt. On the other hand, attention gating utilizes an attention layer (Vaswani et al. 2017), treating learnable prompts as a query and timestep embeddings as key and value, resulting in weighted prompts directly. Upon comparing the two gating architectures, we found that linear gating achieves better FID (5.87) compared to attention gating (6.41). As a result, we adopt linear gating as our default setting.
Figure 5: Prompt activation. Brighter indicates stronger.
(a) Unconditional Image Generation on FFHQ (Karras, Laine, and Aila 2019)
Baseline
+ PT
+ DMP
(b) Text-to-Image Generation on MS-COCO (Lin et al. 2014).
Baseline
+ PT
+ DMP
Figure 6: Qualitative comparison among the baseline (DiT-B/2 (Peebles and Xie 2022)), naive prompt tuning (PT) applied to the baseline, and DMP applied to the baseline on (a) FFHQ and (b) MS-COCO datasets.
@@
Gating type. In Table 4(b), we evaluate two design choices for creating mixture-of-prompts: hard vs. soft selection. With hard selection, we choose the top-$192$ prompts out of 256 prompts (75% of total prompts) based on the gating probabilities, using them only with a weight of 1. Whereas, soft selection uses all prompts but assigns different weights to each. Soft selection leads to further improvement, whereas hard selection results in worse performance. Therefore, we set the soft selection as our default setting.
@@
Prompt selection. By default, DMP inserts learnable prompts into the input space of every blocks in the diffusion model. Two choices arise in this context: (1) uniform: gating function $\mathcal{G}$ in Eq. 3 receives only the timestep embedding ${\bm{t}}$ as input and applies common weights to the prompts at every block, thus prompt selection is consistent across all blocks. (2) distinct: $\mathcal{G}$ processes not only ${\bm{t}}$ but also current block depth $i$ as inputs, generating different weights for each block. As shown in Table 4(c), using distinct prompt selections leads to enhanced performance. Therefore, we input both the timestep embedding and current block depth information to the gating function, enabling the distinct prompt combinations for each block depth as our default setting.
@@
Prompt position. While previous prompt tuning approaches (Li and Liang 2021; Lester, Al-Rfou, and Constant 2021; Jia et al. 2022; Zhou et al. 2022b; Wang et al. 2022) typically prepend learnable prompts to image tokens, we directly add prompts element-wise to image tokens to maintain the input sequence length. Table 4(d) ablates different choices for inserting prompts into the input space and their impact on performance. We compare two methods: prepend vs. add. For “add”, we use 256 prompts to match the number of image tokens, and for “prepend”, we utilize 50 prompts. Although ”prepend” should ideally use 256 tokens for a fair comparison, we limit it to 50 tokens due to severe divergence, even when both methods are equally zero-initialized. The results show that “add” method improves performance with a stable optimization process, achieving a 5.87 FID compared to the baseline of 6.27 FID. On the other hand, the “prepend” leads to a drop in performance, with an FID of 6.79. Additionally, “add” has the advantage of not increasing computation overhead. Based on these findings, we set “add” as our default for stable optimization.